OpenStudy (anonymous):

Part A- Explain why the xcoordinates of the points where the graphs of the equations y=4^x and y=2^x^+^2 intersect are the solutions of the equation 4^x = 2^x^+^2?

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

marry me? yes no? :DDD

OpenStudy (anonymous):

uhm nothing special here..... and can you help with this problem?

OpenStudy (campbell_st):

The point of intersection is a common point, which is part of both curves. The x value is common to both and produces the commo y value from the point when substituted into both equations.

OpenStudy (anonymous):

so just find the point of intersection? @campbell_st

OpenStudy (anonymous):

uhmmm

OpenStudy (campbell_st):

well it can be solved aglebraically but you need the same base... \[4^x = (2^2)^x = 2^{2x}\] so it looks like you equation can be written as \[2^{2x} = 2^{x + 2}\] is this is the case then you just equate the powers and solve 2x = x + 2

OpenStudy (campbell_st):

another solution that can be used is to just use a graphing package or site like https://www.desmos.com/calculator

OpenStudy (anonymous):

so wait the point f intersection would be 2x?

OpenStudy (campbell_st):

no if I had your equation correct the x value of the point of intersection is the solution to 2x = x + 2

OpenStudy (campbell_st):

if you solve this x = 2 so to check does \[4^2 = 2^{2 + 2}~~?\]

OpenStudy (anonymous):

oh ok thank you so could you help me with part b and c too?

OpenStudy (campbell_st):

so what is part b and c

OpenStudy (campbell_st):

can I check the equation is \[4^x = 2^{~x + 2}\]

OpenStudy (anonymous):

oh oops lol sorry here Part B- Make tables to find the solution to 4^x = 2^x^+^2. Take the integer values of x between -3 and 3.

OpenStudy (anonymous):

and yes

OpenStudy (campbell_st):

ok... so table of values so substitute the x values into each equation x | -3 | -2 | -1 | 0 | 1 | 2 | 3 | -------------------------------------------- 4^x | 1/64 | 1/16 | 1/4 | 0 | 4 | 16 | 64| -------------------------------------------- 2^(x + 2) | 1/2 | 1 | 2 | 4 | 8 | 16 | 32 and you see when x = 2 4^x = 16 and 2^(x + 2= 16

OpenStudy (anonymous):

ok you los me I know how to substitute them but there are 2 number?.....

OpenStudy (campbell_st):

well rather than do 2 tables I used 1 set of x values and have 2 rows for the different y values 1st row is \[y = 4^x\] 2nd row is \[y = 2^{~x + 2}\]

OpenStudy (campbell_st):

so reading the table, the point of intersection is (2, 16)

OpenStudy (anonymous):

uhm sorry today just isn't my day im having a total brain fart right not ok ok im kinda following

OpenStudy (campbell_st):

I can't help with C as I'm in Australia and need to go to class.

OpenStudy (campbell_st):

hope it all makes some sense